Italian Barolo ~ Italy is famous for several types of wine. The region of Piedmont is famous for a type of wine called “Barolo”. A wine salesperson wants to know the demand for Barolo wine sold at Supermarkets. He selects a random sample of wine sales from a popular supermarket and constructs a 95% confidence interval for the actual proportion of Barolo wine sold to be (0.2672, 0.4259). What is the margin of error? Give the exact answer as a decimal.
Answer #1

Solution :

Point estimate = = (Lower confidence interval + Upper confidence interval ) / 2

Point estimate = = (0.2672+0.4259) / 2=0.34655

Margin of error = E = Upper confidence interval - = 0.4259-0.34655=0.07935